OpenStudy (anonymous):

(2x+3)(x-1)=0..Zero Property.

OpenStudy (anonymous):

OK jjkm18 what are your two numbers in your problem

OpenStudy (anonymous):

3 & -1..?

OpenStudy (anonymous):

The stuff in the parentheses are the numbers 2x + 3 is a number and x - 1 is a number.. OK?

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

So either 2x + 3 = 0 or x - 1 = 0 do you know how to solve for x?

OpenStudy (anonymous):

No. I forqot

OpenStudy (anonymous):

2x + 3 = 0 you need to get the x by itself so you -3 from both sides 2x + 3 - 3 = 0 - 3 which leaves you with 2x = -3 then you would divide both sides by 2 2x = -3 -- -- and you would get 2 2 x = -3/2 then solve x - 1 = 0 you would add 1 to both sides x - 1 + 1 = 0 + 1 x = 1 so your answers are x = -3/2 or x = 1

OpenStudy (anonymous):

Thanks, I KINDOF get it now..